UT Feb 25 2021



Night Partition

  • Night split 50/50 w/PEPSI science - D/E-time 2nd half
  • Night Center 00:33 - handover 00:28
  • Moon 94% effectively up all night

Night Plan

  • AO Engineering Time - highest priority - (LBTO AO and Arcetri)
    • LUCI-SOUL AO tasks - test 6 different PAs per object
    • LUCI-SOUL Arcetri tests
      • closed loop on AO ref stars (on-axis) to collect SR and telmetry
      • one or more stars in the following categories
        • DONE (1 completed) 14.5 < Rmag < 15.5 Fast Bin 3
        • 15.8 < Rmag < 16.5 Slow Bin 3
        • 16.51 < Rmag < 17.0 Fast Bin 4
        • 17.0 < Rmag < 18.0 Slow Bin 4
        • Rmag > 18.0 (as far as we can go....)
    • LUCI-SOUL ESM tasks - 1hr - only if no other full AO
      • Hoag's Object - on-axis AO Ref star (galaxy core as AO Ref star R=13.44 predicted WFSmag =14.7)
  • if cannot switch to LUCI due to M3 SX issue:
    • LBC Solar System objects (lost asteroids) - 2hrs (Veillet)
    • MODS Local Mergers - 1.5hrs per target (Rothberg)
    • MODS Field Correction check for IT 8315
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
      • run IDL task field_collect (same as measuring field aberrations)
        • LDG and RDG in parallel
        • Use 6-8 stars
  • Poor Weather Backup
    • ESM data collection (Bino-LUCIs 1hr per target)
      • M67 visible from handover until 03hr
      • M92 visible from 2am through twilight
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • MODS Local Mergers (Bino-MODS 90min per target based on conditions) - Rothberg
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
    • ICRF3 (bino-LUCIs) - Rothberg
      • ESM has priority over ICRF3 - if similar conditions collect ESM data
      • /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/ICRF3

Overview:

All times are in UT:

20210224 22:20 Reboot of obs2-4

20210224 23:40 We want to set the SX M3 early to ensure the selector does not have issues. Geno is going ahead for us (Thanks Geno!).
Geno has set the M3 to the PEPSI PFU without issue. Geno mentioned somethign about the right M1 was in a limit, but he was going to pass that on to David.

20210224 23:47 Powering up the SX WFS for later. w_stop_all/start_all, processes all connected first try to message_D. Power up good first try. Leaving at ambient until closer to handover time.

20210225 0:00 Verify field stop alignment. Noticed some crud in N30 images. Detector was out of focus. Switching to N375 resolved detector focus issue and verified that the field lens was clear.

20210225 0:50 Taking some luci darks for the AO.
  • luci1.20210225.0015-24 1x2.5
  • luci1.20210225.0025-29 10x2.5
  • luci1.20210225.0030-34 26x2.5
  • luci1.20210225.0035-39 2x15.0
  • luci1.20210225.0040-44 2x12.5
  • luci1.20210225.0045-49 2x8.5
  • luci1.20210225.0050-54 2x4.5
20210225 1:17 Sunset

20210225 1:21 Opening

20210225 1:38 Pointing and collimation on a nearby star and then we will move on to our first PEPSI target of the night: V501 Aur.

20210225 2:04 Science started on V501 Aur

20210225 2:07 12 deg evening twilight

20210225 2:11 Winds are too high. TSS enabled. We need to move. Aborting. David recovering the adsec

20210225 2:29 Going to next target. This one is a little better but we may end up with the same issue before it is complete. HD45762 (V8.24). Pointing correction needed.
100micron fiber both, BS1 CD2 CD4 first then CD1 &CD6

20210225 3:39 preset to V501Aur, the position is better releative to the moon. We saw some weird behaviour on the right guider with it oscillating up and down in and out of the box. Resending the preset resolved.
200 micron fiber Ag mirror both, BS2 CD3 CD5 first then CD2 &CD4. Seeing is 0.5-0.6"

20210225 5:18 jitter on left causing guide star to jump out of the field for the last minute or so of the exposure. We are moving on to the next target anyways but worth noting. Going to 2MASS J09290939-08. Need a pointing check again.

20210225 5:43 Wind gusts over limit. Closing

20210225 7:21 Handover between PEPSI and D, reconfigure to LUCI

20210225 7:32 Night Center

20210225 7:37 We have reopened! Authorized for LUCI1. Pointing check and collimation

20210225 Bin3 Test:
R=14.9 (expected WFS is 14.9) K=10.5. seeing ~1"
PA=0 el=80, rotator=262 camera lens saturated
PA=60 el=77, rotator=323. the moon is causing issues with the DIMM. Moving to another field

R=14.3 K11.4 (LHS 2763)
Pointing was off again requiring another pointing check
PA=0, el=59, rotator 441.6 . Enrico and team are monitoring DIMM. WFS magnitude 15.2, seeing ~1.4". Bin3, frequency 826, 3.8 photons/subapp, ~20 average pupil intensity,
Enrico and team are taking data with luci and some telemetry
taking luci and telemetry. Seeing deteriorating. 1.7", rinse and repeat on this target (same PA and preset, just to checkout the params)
WFS seeing 15.1 estimated magnitude this time. Bin3, frequency 811, 3.9 photons/subapp, ~20 average pupil intensity,
taking luci data and telemetry
offset 20" in X and 20" in Y: sky frames 0064. Camera lens definitely moved when we offset to sky position.
Safeskipped, but reason unknown. GOpt was stable. Seeing 1.1"

20210225 9:26 Preset to next bin4 target:
R=15, K=13.7 PA=0 el = 82 Rotator 149.5 (1291-0230904)
OCAM illumination error when we went. Pointing off again after only a 10 degree move!
Bin4 frequency 471. Optimize gain is failing, putting some very high HO1 on there.
Safeskip and lost target. too much light from the clouds.

2021022510:06-10:31 Idle for clouds.
Resent preset to the SAME TARGET we were tracking, and some how the GS and AOrefer had drifted so much that we need another pointing check.
R=15, K=13.7 PA=0 el = 80 Rotator 240 (1291-0230904). Estimated mag 16.4. Bin3, freq 269. 3.4 photons/subapp. 18 average pupil intensity. failed optimize gain. the sky is pretty cloudy.

20210225 10:55 Moving on to brighter targets because of the clouds. R=12.7 K=8.2 (G63-50). Shocking, pointing was off again and required a pointing check.

PA 0 el 65 rotator 354 camera lens saturated bin 2 frequency 200 sent
PA 60 el 65 rotator 412, WFS estimated magnitude is 13.6, closed with bin2 frequency 886. taking luci data and telemetry
possible binary?

20210225 11:37 thick clouds now. taking some P-Beta to go with out Brackett Gam to investigate teh structure while we wait for these clouds to thin.

Enrico is experiencing an issue with his routine that he uses for collecting tracking data. Investigating.
  • Fabio is cycling wfs arbitrator
  • Reclosed on the same target. clouds have lifted, WFS estimated mag is 12.6. loop is bin 1 freq 718
20210225 12:19 Clouds have cleared out for the most part. We are going to move back to R=15 K=13.7 (1291-0230904). Pointing off again. Pointing check need. Nothing after first pointing check. We needed a second one.

PA=0 el 59 rotator 277 camera lens saturated (bin 3)
Sending for PA=60: rotator 338, el 58: bin 3 at 253 Hz . Seeing variable, up to 1.45". Safe skipping when optimizing. Greg manually applying gains. Estimated mag on WFS is 16.6, total incoming light 2.9 photons/subapp, average pupil intensity ~14.
Seeing deteriorating, up to 1.9 while greg tried to apply optical gain... second try worked though.

20210225 12:54 we are 2"+ seeing. this is not useful. Calling it a night

20210225 12:58 12 deg morning twilight

20210225 13:48 Sunrise

-- BarryRothberg - 22 Feb 2021
Topic revision: r12 - 25 Feb 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback